# Complete transparency with three active-passive-coupled optical resonators

### v2

@@ Abstract
The phenomena of induced transparency, with the typical examples of electromagnetically induced transparency (EIT) in atomic media and coupled optical resonators, have attracted tremendous interest since their discoveries. Due to the limitations of the involved elements, however, near-100% transmissions were reported under highly demanding experimental conditions for atomic and other media. Based on a structure of three linearly coupled optical resonators, an active one carrying a possibly arbitrary optical gain and two passive ones simply with dissipation, we demonstrate that a transmitted light field can become completely transparent through the structure, which displays all properties similar to those of EIT. A destructive interference mechanism is discovered to totally eliminate the intracavity field in the resonator directly coupled to the transmitted field of any feasible power, when the coupling strength of two other resonators is tuned to a point inside the designed range. This mechanism works for all possible coupling strengths of the dark resonator with its neighboring resonator and the input field, as well as for any available quality factor from its fabrication. The transparency window size and the output field intensity at the transparency window center can be freely engineered by adjusting two inter-cavity couplings, without modifying the built-in optical gain which can be just slightly stronger than the dissipation of the active resonator.
@@
with $\Delta_{i}=\omega_{p}-\omega_{i}$ ($i=1,2,3$) and $\eta=J_{2}^{2}(\kappa_{1}-i\Delta_{1})$. It can be seen from Eq. (4) that $A_{2}$ and $A_{3}$ cannot be zero because $J_{1}$, $J_{2}$ and $\kappa_{3}$ are non-zero, but $A_{1}$ can be zero under certain conditions. These facts are crucial for achieving a complete transparency. In order to obtain the conditions, we rewrite $A_{1}$ in the form of continued fraction:
@@
Now we consider three resonators with the same resonance frequency $\omega_{1}=\omega_{2}=\omega_{3}=\omega_{0}$, so that Eq. (7) can be simplified to
@@
Instead of being a limit point that should be asymptotically approached, the completely vanishing field in cavity $1$ emerges at a midpoint of tuning the coupling $J_{2}$. As represented by the samples in Fig. 4(a), tuning the coupling strength from $J_{2}>J_{2,c}=\sqrt{\kappa_{2}\kappa_{3}}$ to $J_{2}<J_{2,c}$ will continuously switch the transparency window center from less than the $100\%$ transmission to more than the $100\%$ transmission, while the dispersion property demonstrating the slow-light effect is still preserved as seen from Fig. 4(b). As the distance between cavity $2$ and $3$ keeps increasing from where cavity $1$ is in the vacuum state of completely eliminated intracavity field, which is at $J_{2,c}=0.6$ MHz for the system in Fig. 4, more photons entering cavity $1$ realize an intracavity field phase to strengthen the output, so that a higher and higher ratio $|a_{out}/a_{in}|^{2}>1$ will be achieved until the system loses its dynamical stability at $J_{2}=0.2449$ MHz. Then it becomes something like two coupled resonators in a PT-symmetry broken regime due to the choice $J_{1}<\sqrt{\kappa_{1}\kappa_{2}}$. There is no restriction on the qualities of used optical resonators for seeing these phenomena of amplifying and reducing the output without modifying the built-in optical gain, since the system stability under an arbitrary net gain rate $\kappa_{2}$, which is detailed in Appendix D, and an available coupling strength range covering the point $J_{2,c}$ guarantee their occurrence. The structure in Fig. 1(a) can thus beat a common notion that the better the used optical resonators are, the better the induced transparency will be.
@@
In summary, we have presented an interference mechanism of genuinely reducing the susceptibility $\varepsilon_{T}$ of the structure in Fig. 1(a), which is exactly solvable, to zero under the implementable conditions, so that the realized 100% transmission exists for any feasible power of transmitted fields. Apart from scanning the input frequency, we only require the tuning of two inter-cavity couplings to see its manifestation: setting one inter-cavity coupling at $J_{2,c}=\sqrt{\kappa_{2}\kappa_{3}}$ and independently adjusting the other coupling $J_{1}$ (possibly together with the external coupling $\kappa_{ex}$) for a suitable transparency window size. Rather than being used to cancel the transmission loss, the optical gain creates the point of perfect interference in the first resonator, together with two other parameters $J_{2}$ and $\kappa_{3}$ that are also totally independent of the properties of cavity $1$, and can modify the dispersion of a transmitted field at the same time. The significance of the scenario lies in two aspects. First, to the research of induced transparency based on optical resonator [22, 23], it proves that the ultimate goals of complete transparency and free control on the output/input ratio can be realized without complex designs or particular nonlinear elements. Second, it will non-trivially extend the scope of non-Hermitian photonics based on optical gain. Optical gain in one of two coupled optical resonators provides a realization of PT symmetry [47, 48, 49, 50, 51, 52, 60, 69], and the study of the associated exceptional points has been developed into a direction of intensive research [70, 71, 72]. Coupling one more resonator will have an even more non-trivial scenario involving a special dark mode with its rich properties, and it presents a rare example of something unusual simply built on the common elements like active, passive resonators and their linear couplings.
@@ Appendix A Appendix
These stable regimes are exemplified in Fig. S-2(a). Here we assume that only the two coupling strengths $J_{1}$ and $J_{2}$ are tunable by the relative distances between the resonators, with the practical examples for microspheres [15] and microtoroid resonators [67] (the range of tuning $J_{1}$ and $J_{2}$ of this type of microresonators can be over $1$ GHz), while all other system parameters are fixed after the resonators are fabricated. Fig. 4 in the main text presents a process of tuning the coupling $J_{2}$ while the value of $J_{1}$ is fixed on the left side of the vertical line of $J_{1}=\sqrt{\kappa_{1}\kappa_{2}}$ in Fig. S-2(a). The coupling $J_{2}$ can be continuously tuned downwards across the horizontal line of $J_{2}=\sqrt{\kappa_{2}\kappa_{3}}$ in Fig. S-2(a), on which there is a complete cancellation of the field in cavity $1$, outputting the field from lower than the $100\%$ transmission to higher than the $100\%$ transmission. After the coupling $J_{2}$ touches the lower boundary, the system will enter the regime of dynamical instability, where the optical gain saturation will play a significant role so that the finally stabilized gain rate will be much lower than the original $\kappa_{2,0}$. On the right side of the vertical line of $J_{1}=\sqrt{\kappa_{1}\kappa_{2}}$, the output field intensity at the transparency window center will tend to infinitely high as the coupling $J_{2}$ is approaching to zero, but the associate transparency window will be vanishing at the same time.

### v3

@@ Abstract
The phenomena of induced transparency, with the typical examples of electromagnetically induced transparency (EIT) in atomic media and those based on coupled optical resonators, have attracted tremendous interest since their discoveries. Owing to the limitations of the involved physical elements, however, near-100% transmissions were reported under highly demanding experimental conditions. With a structure of three linearly coupled optical resonators, an active one carrying optical gain and two passive ones simply with dissipation, we demonstrate that a transmitted light field can become completely transparent through the structure, which displays all properties similar to those of EIT. It is due to a destructive interference mechanism that totally eliminates the intracavity field in the dissipative resonator directly coupled to the transmitted field of any feasible power, when the coupling strength of two other resonators is tuned across a point determined by their associated gain and loss rates. This mechanism works for all possible coupling strengths of the dark resonator with the input field and its neighboring resonator, as well as for any available quality factor from its fabrication. The transparency window size and output field intensity can be freely adjusted by tuning two inter-cavity couplings of wide ranges, without modifying the built-in optical gain which can be just slightly stronger than the dissipation of the active resonator.
@@
with $\Delta_{i}=\omega_{p}-\omega_{i}$ ($i=1,2,3$) and $\eta=J_{2}^{2}(\kappa_{1}-i\Delta_{1})$. $A_{2}$ and $A_{3}$ cannot be zero because $J_{1}$, $J_{2}$ and $\kappa_{3}$ are non-zero, but $A_{1}$ can be zero under certain conditions. This is a unique feature of the structure in Fig. 1(a) and is crucial for achieving a complete transparency. To see the conditions for complete transparency more clearly, we rewrite $A_{1}$ in the form of continued fraction:
@@
Now we consider three microresonators with the same resonance frequency $\omega_{1}=\omega_{2}=\omega_{3}=\omega_{0}$ from their fabrication and the control of their temperatures [60], so that Eq. (7) will be simplified to
@@
Instead of being a limit point that should be asymptotically approached, the completely vanishing field in cavity $1$ emerges at a midpoint of tuning the coupling $J_{2}$. As represented by the samples in Fig. 4(a), tuning the coupling strength from $J_{2}>J_{2,c}=\sqrt{\kappa_{2}\kappa_{3}}$ to $J_{2}<J_{2,c}$ will continuously switch the transparency window center from less than the $100\%$ transmission to more than the $100\%$ transmission, while the dispersion property demonstrating the slow-light effect is still preserved as seen from Fig. 4(b). As the distance between cavity $2$ and $3$ keeps increasing from where cavity $1$ is in the vacuum state of completely eliminated intracavity field, which is at $J_{2,c}=0.6$ MHz for the system in Fig. 4, more and more photons will enter cavity $1$ and realize an intracavity field phase for the mode $a_{1}$ to interfere with the input field constructively, so that a higher and higher ratio $|a_{out}/a_{in}|^{2}>1$ will be achieved until the system loses its dynamical stability at $J_{2}=0.2449$ MHz. Then it becomes something like two coupled resonators in a PT-symmetry broken regime due to the choice $J_{1}<\sqrt{\kappa_{1}\kappa_{2}}$. There is no restriction on the qualities of used optical resonators for seeing these phenomena of amplifying and reducing the output without modifying the built-in optical gain, since the system stability under an arbitrary net gain rate $\kappa_{2}$, which is detailed in Appendix D, and an available coupling strength range covering the point $J_{2,c}$ guarantee their occurrence. The structure in Fig. 1(a) can thus beat a common notion that the better the used optical resonators are, the better the induced transparency will be.
@@
In summary, we have detailed an interference mechanism that genuinely reduces the susceptibility $\varepsilon_{T}$ of the structure in Fig. 1(a) to zero under the implementable conditions, so that a light field transmitted at the resonance frequency of those coupled resonators and with any feasible power will become totally invisible to the structure, no matter how strongly it is coupled to the first resonator. Besides scanning the transmitted field frequency, we only require the adjustments of two inter-cavity couplings to see its manifestation: setting one inter-cavity coupling at $J_{2,c}=\sqrt{\kappa_{2}\kappa_{3}}$, and independently choosing the other coupling $J_{1}$ for a suitable transparency window size. Rather than being used to cancel the transmission loss, the optical gain with its arbitrary net gain rate $\kappa_{2}$ creates the point of perfect interference in the first resonator, together with two other parameters $J_{2}$ and $\kappa_{3}$ that are also totally independent of the properties of cavity $1$, and can modify the dispersion of a transmitted field with the varied $\kappa_{2}$. The significance of the scenario lies in two aspects. First, to the research of induced transparency based on optical resonator [22, 23], it proves that the ultimate goals of complete transparency and free control on the output/input ratio can be realized without complex designs or particular nonlinear elements. Second, it will non-trivially extend the scope of non-Hermitian photonics based on optical gain. Optical gain in one of two coupled optical resonators provides a realization of PT symmetry [47, 48, 49, 50, 51, 52, 60, 69], and the study of the associated exceptional points has been developed into a direction of intensive research [70, 71, 72]. Coupling one more resonator will have an even more non-trivial scenario involving a special dark mode with its rich properties, and it presents a rare example of something unusual simply built on the common elements like active, passive resonators and their linear couplings.
@@ Appendix A Appendix
These stable regimes distribute as those exemplified in Fig. S-2(a). Here we assume that only the two coupling strengths $J_{1}$ and $J_{2}$ are tunable by the relative distances between the resonators, with the practical examples for microspheres [15] and microtoroid resonators [67] (the range of tuning $J_{1}$ and $J_{2}$ of this type of microresonators can be over $1$ GHz), while all other system parameters are fixed after the resonators are fabricated. Fig. 4 in the main text presents a process of tuning the coupling $J_{2}$ while the value of $J_{1}$ is fixed on the left side of the vertical line of $J_{1}=\sqrt{\kappa_{1}\kappa_{2}}$ in Fig. S-2(a). The coupling $J_{2}$ can be continuously tuned downwards across the horizontal line of $J_{2}=\sqrt{\kappa_{2}\kappa_{3}}$ in Fig. S-2(a), on which there is a complete cancellation of the field in cavity $1$, outputting the field from lower than the $100\%$ transmission to higher than the $100\%$ transmission. After the coupling $J_{2}$ touches the lower boundary, the system will enter the regime of dynamical instability, where the optical gain saturation will play a significant role so that the finally stabilized gain rate will be much lower than the original $\kappa_{2,0}$. On the vertical line of $J_{1}=\sqrt{\kappa_{1}\kappa_{2}}$, the output field intensity at the transparency window center will tend to infinitely high as the coupling $J_{2}$ is approaching to zero, but the associate transparency window will be vanishing at the same time. The output will always tend to infinity if the system approaches the boundary between the stable and unstable regime, except for the point $(J_{1},J_{2})=(0,\sqrt{\kappa_{2}\kappa_{3}})$. Going to the right side of the vertical line $J_{1}=\sqrt{\kappa_{1}\kappa_{2}}$, the system will behave like two coupled resonators in a PT-symmetric regime when one of the couplings is reduced to $J_{2}=0$, and the output with a fixed $J_{1}$ reaches the maximum there. It is clearly seen from Fig. S-2(a) that the smaller the net gain rate $\kappa_{2}$ is (with the fixed other built-in parameters), the smaller the unstable regime of the system will be shrunk to.